# Bidimensional measurements of photon statistics within a multimodal temporal framework

### v1

@@ I Introduction
For decades, researchers have sought the ability to faithfully image spatial bi-dimensional information generated at incredibly fast timescales. Ultrafast chaotic behaviors, complex nonlinear optical phenomena, and rapid biological processes like protein rearrangement are just a few examples of events that occur on timescales ranging from picoseconds to femtoseconds. Additionally, transient events, specific to two-dimensional systems, reflect many major fundamental mechanisms occurring in biology, physics, or chemistry [Bard_2001_book, Gorkhover2016, Imada1998].
Traditionally, pump-probe methods [Hockett2011, Wong2012, Acremann2000, Feurer2003, Kodama2001, Velten2012, Barty2008, Hajdu2000, Zewail1988] have been used to capture these fast dynamics through repeated measurements. However, many ultrafast phenomena are either non-repetitive or challenging to reproduce [Jalali2010, Solli2007, Poulin2006, Tuchin2007, iaulys2014, Sciamanna2015]. Significant efforts have been made to create platforms capable of performing ultrafast single-shot optical measurements [Liang2018, Gao2014, Jin2024], intending to be free of the paradigmatic repeatable measurement schemes. The main objective of these platforms is the possibility of measuring bi-dimensional photonic fields, at an ultrafast scale (picosecond to femtosecond information with high repetition rate) without alteration of the information of interest. Well-established strategies consist of using cutting-edge pulsed laser technologies to probe a system with strong optical pulses to have enough signal-to-noise ratio in a single-shot experiment and then retrieving the information of interest with various elaborated methods [Wang2014, Abramson1978, Li2014, Nakagawa2014, Yue2017, Ehn2017]
@@ III.1 Singular Value Decomposition, Bogoliubov transformation and input-ouput relation
This part introduces the mathematical tools and conventions to perform the modal treatment of the DFG process and follows references  [Wasilewski2006, Patera2012, Christ2013]. We decompose the positive part of the electric field in terms of the Fourier components
@@
$\delta(\Omega-\Omega^{\prime})$ and $\delta(z-z^{\prime})$ are Dirac delta functions and $\delta_{n,n^{\prime}}$ is the Kronecker delta. The index $n=\{\mathrm{s},\mathrm{i}\}$ accounts for the signal and idler fields collinearly propagating along the $z$-axis through the nonlinear medium. The interactions between these fields are mediated by a strong classical pump pulse of spectral amplitude $A_{\mathrm{p}}$, and is such that we have perfect phase matching at carrier frequencies: $\omega_{\mathrm{p}}=\omega_{\mathrm{s}}+\omega_{\mathrm{i}}$ and $\bm{k}_{\mathrm{p}}=\bm{k}_{\mathrm{s}}+\bm{k}_{\mathrm{i}}$. The wavevectors $\bm{k}_{\mathrm{p}}$, $\bm{k}_{\mathrm{s}}$, and $\bm{k}_{\mathrm{i}}$ are associated to the pump, signal, and idler fields, respectively. In the undepleted pump approximation (the pump amplitude does not vary over the propagation through the crystal i.e., it does not depend on $z$), the evolution of signal and idler is given by [Christ2013]
@@
Since Eqs. (7) and (8) are linear in the field operators, a formal solution can be obtained in terms of a linear integral transformation (see for example [Wasilewski2006, Patera2012, Christ2013]), generalizing the usual Bogoliubov transformation. Since $K$ is $z$-dependent, the exact solution can only be obtained by numerical integration. On the other hand, it has been shown [Christ2013] that the Magnus expansion truncated at first order provides an accurate analytical description up to a degree of amplification of about 1212dB in fields, which corresponds to an amplification in photon number of a factor 16. In other words, this approximation is valid as long as the system gain is smaller than 16. The advantage of using truncated Magnus expansion, as we will show below, is that the solution can be cast in explicit form in terms of the normal modes of the evolution, offering a visualization and understanding of the modes in which photons are emitted. These modes are also known as “supermodes” or “temporal” modes; we will use the denomination “temporal modes” in this paper.
@@
where $\psi_{m}(\Omega)$ and $\phi_{m}(\Omega^{\prime})$ are the temporal eigenmodes of the signal and the idler, respectively, and $\lambda_{m}\geq 0$ is the corresponding singular value. It is interesting to note that this decomposition also induces a SVD of Eq. (12) that preserves the symplectic structure of the problem, and it is known as Bloch-Messiah decomposition [Bennink2002, Law2000, deValcrcel2006, Wasilewski2006, Patera2012, Christ2013]. The latter can be written as
@@ III.2 Analytical and numerical structure of the temporal modes
which is plotted in Fig. 5(b). It shows the strength of the frequency coupling in the nonlinear crystal, accounting for the spectral limitation of the pump. The pump’s characterization is provided in the appendix. As explained in the previous part, temporal modes can generally be obtained through the numerical SVD of the JSA. However, apart from configurations where the pump and phase-matching are specifically engineered to tailor quantum states [Patera2012, Serino2024], the JSA in most experimental situations can be well approximated by a bi-variate that we define as $M_{1}^{\text{ana}}(\Omega_{\mathrm{i}},\Omega_{\mathrm{s}})$:
@@ III.3 Input-Output formalism and vacuum statistics
From Eq. (LABEL:eq:SM_Input_Output), it is possible to compute the probability of having $n$ photons in the idler mode $m$ after the crystal (experimental situation of interest), starting from the vacuum. For a given $m$, we can call $|0\rangle_{\rm in}$ the vacuum of the input modes (idler and signal), such that $c^{m}_{\mathrm{s}}|0\rangle_{\rm in}=c^{m}_{\mathrm{i}}|0\rangle_{\rm in}=0$. It is convenient to introduce the state $|0\rangle_{\rm out}$, the vacuum of the $b$ operators, $b^{m}_{\mathrm{i}}|0\rangle_{\rm out}=b^{m}_{\mathrm{s}}|0\rangle_{\rm out}=0$. Following [Takagi1986], one easily finds that
@@ III.5 Calculation of 0D Photon statistics with input coherent state
As the temporal modes are independent, the calculation is performed for a single set of modes $m$, thus, the index $m$ is suppressed for clarity. We first compute the distribution of the number of output photons in the idler when starting from $n_{0}$ input signal photons. Similar results have been obtained in the context of Hawking radiation [Parker1975, Gasperini1990]. We start by computing
@@ V CONCLUSION AND OUTLOOK
Beyond this proof of principle, our approach establishes a robust optical platform for bidimensional photon statistics measurements. Such a capability is particularly relevant for hybrid light–matter systems, where photon correlations encode key signatures of many-body physics. In exciton–polariton microcavities, for instance, single-shot access to photon statistics in two dimensions could provide crucial insights into the onset of superfluidity [Carusotto2013, Amo2009, Amo2009b, Amo2011, Sanvitto2010, Fontaine2022, Caputo2017] , the role of fluctuations, and the emergence of turbulent behavior[Claude2020, Amelio2020] driven by vortex interactions.
Looking ahead, several promising directions emerge from this work. First, the integration of engineered nonlinear materials, such as periodically poled crystals, could offer increased flexibility in phase-matching and allow tailoring of the temporal mode structure. Considering Type-II phase-matching could offer twofold benefits: the possibility of separating signal and idler based on polarization, and the opportunity to engineer the process to operate in a temporally single-mode regime. Finally, other nonlinear processes could also be considered for one-shot two-dimensional photon statistics measurements, such as sum-frequency generation or four-wave mixing. The use of SFG could allow direct access to the input statistics, without the need to infer them as we have done here, provided that the $100\%$ efficiency limit is reached. While in single-pass configurations, such as those in [Eckstein2011, Christ2013, Reddy2013], the efficiency cannot exceed $80\%$, temporal-mode interferometry schemes, such as those proposed by [Reddy2014], could allow approaching near-unity efficiency.
@@ ACKNOWLEDGMENTS
The data that support the findings of this article are openly available [DataAv].
@@
This part provides a details about the creation of a complex input image using an SLM. The A input single image used in the second part of the manuscript to demonstrate faithful imaging was created by tailoring the coherent input beam from the Ti:Sapph laser with a phase-only Spatial Light Modulator (HOLOEYE Pluto 2.1 LCOS). Doing so we created a complex shape possessing a rich transverse $k_{x}$ and $k_{y}$ distribution. To realize this input state, we developed a homemade MRAF algorithm[MRAF_pasienski08] allowing us to obtain a desired image with $20\%$ overall error. The explicit Python code developed for creating the phase mask can be found in the corresponding GitHub repository [ouahrouche_SLM]. The target intensity, the phase mask applied to the SLM, as well as the recorded image measured with a regular camera at the Fourier plane of the SLM, are represented in Fig. 15.
@@
with $l_{c}$ the crystal length over the propagation axis and $g$ is linked to the medium and fields through [LoudonBook]

### v2

@@ I Introduction
For decades, researchers have sought the ability to faithfully image spatial bi-dimensional information generated at incredibly fast timescales. Ultrafast chaotic behaviors, complex nonlinear optical phenomena, and rapid biological processes like protein rearrangement are just a few examples of events that occur on timescales ranging from picoseconds to femtoseconds. Additionally, transient events, specific to two-dimensional systems, reflect many major fundamental mechanisms occurring in biology, physics, or chemistry [8, 23, 26].
Traditionally, pump-probe methods [25, 53, 3, 19, 29, 50, 9, 24, 55] have been used to capture these fast dynamics through repeated measurements. However, many ultrafast phenomena are either non-repetitive or challenging to reproduce [27, 46, 39, 48, 45, 43]. Significant efforts have been made to create platforms capable of performing ultrafast single-shot optical measurements [32, 21, 28], intending to be free of the paradigmatic repeatable measurement schemes. The main objective of these platforms is the possibility of measuring bi-dimensional photonic fields, at an ultrafast scale (picosecond to femtosecond information with high repetition rate) without alteration of the information of interest. Well-established strategies consist of using cutting-edge pulsed laser technologies to probe a system with strong optical pulses to have enough signal-to-noise ratio in a single-shot experiment and then retrieving the information of interest with various elaborated methods [51, 2, 31, 34, 54, 18]
@@ III.1 Singular Value Decomposition, Bogoliubov transformation and input-ouput relation
This part introduces the mathematical tools and conventions to perform the modal treatment of the DFG process and follows references  [52, 38, 13]. We decompose the positive part of the electric field in terms of the Fourier components
@@
$\delta(\Omega-\Omega^{\prime})$ and $\delta(z-z^{\prime})$ are Dirac delta functions and $\delta_{n,n^{\prime}}$ is the Kronecker delta. The index $n=\{\mathrm{s},\mathrm{i}\}$ accounts for the signal and idler fields collinearly propagating along the $z$-axis through the nonlinear medium. The interactions between these fields are mediated by a strong classical pump pulse of spectral amplitude $A_{\mathrm{p}}$, and is such that we have perfect phase matching at carrier frequencies: $\omega_{\mathrm{p}}=\omega_{\mathrm{s}}+\omega_{\mathrm{i}}$ and $\bm{k}_{\mathrm{p}}=\bm{k}_{\mathrm{s}}+\bm{k}_{\mathrm{i}}$. The wavevectors $\bm{k}_{\mathrm{p}}$, $\bm{k}_{\mathrm{s}}$, and $\bm{k}_{\mathrm{i}}$ are associated to the pump, signal, and idler fields, respectively. In the undepleted pump approximation (the pump amplitude does not vary over the propagation through the crystal i.e., it does not depend on $z$), the evolution of signal and idler is given by [13]
@@
Since Eqs. (7) and (8) are linear in the field operators, a formal solution can be obtained in terms of a linear integral transformation (see for example [52, 38, 13]), generalizing the usual Bogoliubov transformation. Since $K$ is $z$-dependent, the exact solution can only be obtained by numerical integration. On the other hand, it has been shown [13] that the Magnus expansion truncated at first order provides an accurate analytical description up to a degree of amplification of about 1212dB in fields, which corresponds to an amplification in photon number of a factor 16. In other words, this approximation is valid as long as the system gain is smaller than 16. The advantage of using truncated Magnus expansion, as we will show below, is that the solution can be cast in explicit form in terms of the normal modes of the evolution, offering a visualization and understanding of the modes in which photons are emitted. These modes are also known as “supermodes” or “temporal” modes; we will use the denomination “temporal modes” in this paper.
@@
where $\psi_{m}(\Omega)$ and $\phi_{m}(\Omega^{\prime})$ are the temporal eigenmodes of the signal and the idler, respectively, and $\lambda_{m}\geq 0$ is the corresponding singular value. It is interesting to note that this decomposition also induces a SVD of Eq. (12) that preserves the symplectic structure of the problem, and it is known as Bloch-Messiah decomposition [10, 30, 15, 52, 38, 13]. The latter can be written as
@@ III.2 Analytical and numerical structure of the temporal modes
which is plotted in Fig. 5(b). It shows the strength of the frequency coupling in the nonlinear crystal, accounting for the spectral limitation of the pump. The pump’s characterization is provided in the appendix. As explained in the previous part, temporal modes can generally be obtained through the numerical SVD of the JSA. However, apart from configurations where the pump and phase-matching are specifically engineered to tailor quantum states [38, 44], the JSA in most experimental situations can be well approximated by a bi-variate that we define as $M_{1}^{\text{ana}}(\Omega_{\mathrm{i}},\Omega_{\mathrm{s}})$:
@@ III.3 Input-Output formalism and vacuum statistics
From Eq. (LABEL:eq:SM_Input_Output), it is possible to compute the probability of having $n$ photons in the idler mode $m$ after the crystal (experimental situation of interest), starting from the vacuum. For a given $m$, we can call $|0\rangle_{\rm in}$ the vacuum of the input modes (idler and signal), such that $c^{m}_{\mathrm{s}}|0\rangle_{\rm in}=c^{m}_{\mathrm{i}}|0\rangle_{\rm in}=0$. It is convenient to introduce the state $|0\rangle_{\rm out}$, the vacuum of the $b$ operators, $b^{m}_{\mathrm{i}}|0\rangle_{\rm out}=b^{m}_{\mathrm{s}}|0\rangle_{\rm out}=0$. Following [47], one easily finds that
@@ III.5 Calculation of 0D Photon statistics with input coherent state
As the temporal modes are independent, the calculation is performed for a single set of modes $m$, thus, the index $m$ is suppressed for clarity. We first compute the distribution of the number of output photons in the idler when starting from $n_{0}$ input signal photons. Similar results have been obtained in the context of Hawking radiation [36, 22]. We start by computing
@@ V CONCLUSION AND OUTLOOK
Beyond this proof of principle, our approach establishes a robust optical platform for bidimensional photon statistics measurements. Such a capability is particularly relevant for hybrid light–matter systems, where photon correlations encode key signatures of many-body physics. In exciton–polariton microcavities, for instance, single-shot access to photon statistics in two dimensions could provide crucial insights into the onset of superfluidity [12, 7, 6, 5, 42, 20, 11] , the role of fluctuations, and the emergence of turbulent behavior[14, 4] driven by vortex interactions.
Looking ahead, several promising directions emerge from this work. First, the integration of engineered nonlinear materials, such as periodically poled crystals, could offer increased flexibility in phase-matching and allow tailoring of the temporal mode structure. Considering Type-II phase-matching could offer twofold benefits: the possibility of separating signal and idler based on polarization, and the opportunity to engineer the process to operate in a temporally single-mode regime. Finally, other nonlinear processes could also be considered for one-shot two-dimensional photon statistics measurements, such as sum-frequency generation or four-wave mixing. The use of SFG could allow direct access to the input statistics, without the need to infer them as we have done here, provided that the $100\%$ efficiency limit is reached. While in single-pass configurations, such as those in [17, 13, 40], the efficiency cannot exceed $80\%$, temporal-mode interferometry schemes, such as those proposed by [41], could allow approaching near-unity efficiency.
@@ ACKNOWLEDGMENTS
The data that support the findings of this article are openly available [1].
@@
This part provides a details about the creation of a complex input image using an SLM. The A input single image used in the second part of the manuscript to demonstrate faithful imaging was created by tailoring the coherent input beam from the Ti:Sapph laser with a phase-only Spatial Light Modulator (HOLOEYE Pluto 2.1 LCOS). Doing so we created a complex shape possessing a rich transverse $k_{x}$ and $k_{y}$ distribution. To realize this input state, we developed a homemade MRAF algorithm[37] allowing us to obtain a desired image with $20\%$ overall error. The explicit Python code developed for creating the phase mask can be found in the corresponding GitHub repository [35]. The target intensity, the phase mask applied to the SLM, as well as the recorded image measured with a regular camera at the Fourier plane of the SLM, are represented in Fig. 15.
@@
with $l_{c}$ the crystal length over the propagation axis and $g$ is linked to the medium and fields through [33]